# bibliography.bib

### v3

@@ 7.1 The Bruin–Troubetzkoy family
7.2 Parameter space of $\mathrm{ITM}$’s on $r$ intervals and future work
In this final subsection, we discuss the properties of the full parameter space $\mathrm{ITM}(r)$ of all $\mathrm{ITM}$’s on $r$ intervals, and we present a few conjectures. Recall that a connected component of the set of stable maps $\mathcal{S}(r)$ corresponds to the maximal neighbourhood $[T]$ of a stable map $T$ that satisfies the three properties from the definition of stability 5.1. Such sets $[T]$ are called stable regions. A very interesting problem is to describe the geometric shape of these regions.
Question 7.2.
@@ 7.2 Parameter space of $\mathrm{ITM}$’s on $r$ intervals and future work
The following example shows that these regions are not convex in general:
Figure 10: Dynamics of the map $T$ from Example 7.3.
Example 7.3 (Stable regions are not convex).
Consider the map $T$ with dynamics as shown in Figure 10. Its non-wandering set consists of the orbit of a single interval $J$ containing $\beta_{4}$ such that the return map $R_{J}$ to $J$ has two branches $J_{1}$ and $J_{2}$. Moreover, we assume that:
  1. 1.
    $R_{J}(J_{2})=T(J_{2})$;
  2. 2.
    $R_{J}(J_{1})=T^{2}(J_{1})$;
  3. 3.
    $T(J_{1})$ is contained in the interior of $I_{6}=[\beta_{5},1)$;
  4. 4.
    $T(\beta_{5}^{+})$ is contained in the interior of $I_{3}=[\beta_{2},\beta_{3})$ (this makes $T$ irreducible);
  5. 5.
    $T(I_{2})$ is contained in the interior of $J$;
  6. 6.
    $T(0)$ and $T(\beta^{-}_{3})$ are contained in the interior of $I_{2}$;
  7. 7.
    $\beta_{1}^{-}+\gamma_{1}=\beta_{2}^{-}$;
  8. 8.
    $\beta_{2}^{+}+\gamma_{3}=\beta_{1}^{+}$.
Thus $T$ is a finite type $\mathrm{ITM}$ on $6$ intervals that satisfies A1,A2 and Matching, but it is not stable because it violates A3, due to equations 7. and 8. Intuitively speaking, if we perturb this map in a special direction, we get a periodic interval which is a definite distance from the non-wandering set of $T$, while if we perturb in any other direction the non-wandering set moves homeomorphically. We will formalise this analysis below.
Figure 11: This is a schematic picture of the parameter space $\mathrm{ITM}(6)$ near the map $T$ from Example 7.3. The picture is $3$-dimensional, while the actual space $\mathrm{ITM}(6)$ is $11$-dimensional.
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be the hyperplanes in $\mathrm{ITM}(6)$ consisting of parameters $(\gamma\,\beta)$ that satisfy equations 7. and 8., respectively. Let $\mathcal{H}_{0}$ be the hyperplane such that $\gamma_{1}+\gamma_{3}=0$. Let $\mathcal{G}:=\mathcal{H}_{0}\cap\mathcal{H}_{1}\cap\mathcal{H}_{2}$, and note that $T\in\mathcal{G}$. A schematic picture in the parameter space locally near $T$ is shown in Figure 11. Note that $\mathcal{H}_{1}$ splits $\mathrm{ITM}(6)$ into two regions such that $\beta_{1}^{-}+\gamma_{1}-\beta_{2}^{-}>0$ holds on one of them, and $\beta_{1}^{-}+\gamma_{1}-\beta_{2}^{-}\leq 0$ holds on the other. Analogously, $\mathcal{H}_{2}$ splits $\mathrm{ITM}(6)$ into two regions, on which $\beta_{2}^{+}+\gamma_{3}-\beta_{1}^{+}\geq 0$ and $\beta_{2}^{+}+\gamma_{3}-\beta_{1}^{+}<0$, respectively. Thus we get that $\mathrm{ITM}(6)$ splits into $4$ regions (+,+),(+,),(,+)(+,+),(+,-),(-,+)( + , + ) , ( + , - ) , ( - , + ) and (,)(-,-)( - , - ), depending on the sign on the equations corresponding to $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$. We will show that all sufficiently small perturbations of $X$, except those contained in $(\mathcal{H}_{0}\cap(+,-))\cup\mathcal{G}$ (the shaded region in Figure 11), are stable. Thus they all belong to a single stable region, which shows that this stable region is not convex. Let $\tilde{T}$ be a sufficiently small perturbation of $T$, so that properties 1.-6. still hold for the corresponding continuations. We have $4$ cases:
Case 1 : $\tilde{T}\in(+,+)$ Then $\tilde{\beta}_{1}^{-}+\tilde{\gamma}_{1}=\beta_{2}^{-}+\epsilon_{1}$ and $\tilde{\beta}_{2}^{+}+\tilde{\gamma}_{3}=\tilde{\beta}_{1}^{+}+\epsilon_{2}$, where $\epsilon_{1},\epsilon_{2}>0$ are small. Thus $\tilde{T}^{2}(\tilde{\beta}_{1}^{-})$ and $\tilde{T}^{2}(\tilde{\beta}_{2}^{+})$ are both contained in the interior of $\tilde{J}$, by property 3. Thus $\tilde{T}$ does not violate A3, so it is a stable map.
Case 2 : $\tilde{T}\in(+,-)$ Then $\tilde{\beta}_{1}^{-}+\tilde{\gamma}_{1}=\beta_{2}^{-}-\epsilon_{1}$ and $\tilde{\beta}_{2}^{+}+\tilde{\gamma}_{3}=\tilde{\beta}_{1}^{+}+\epsilon_{2}$, where $\epsilon_{1},\epsilon_{2}>0$. If $\epsilon_{1}=\epsilon_{1}$, then the interval $[\beta_{1}-\epsilon_{1},\beta_{1})$ is periodic with period $2$, so $\tilde{T}$ is not stable. Let us now assume that $\epsilon_{1}>\epsilon_{2}$, with the other case being analogous. Then $\tilde{T}^{2}(\tilde{\beta}_{1}^{-})$ must be contained in $\tilde{I}_{2}$, so $\tilde{T}^{3}(\tilde{\beta}_{1}^{-})$ is contained in the interior of $\tilde{J}$, by property $3$. Thus $\tilde{\beta}_{2}^{-}$ also gets eventually mapped into $\tilde{J}$ by iterates of $\tilde{T}$. Thus $\tilde{T}$ does not violate A3 in this case as well, so it is a stable map.
Case 3 : $\tilde{T}\in(-,-)$ It is clear that in this case we have that $\tilde{T}(\tilde{\beta}_{1}^{-}),\tilde{T}^{2}(\tilde{\beta}_{2}^{+})\in\tilde{I}_{2}$, so $\tilde{T}^{2}(\tilde{\beta}_{1}^{-}),\tilde{T}^{3}(\tilde{\beta}_{2}^{+})\in\tilde{J}$, by property 3. Thus $\tilde{T}$ does not violate A3, so it is stable in this case as well.
Case 4 : $\tilde{T}\in(-,-)$ Similar to Case 3, we have that $\tilde{T}^{3}(\tilde{\beta}_{1}^{-})\in\tilde{J}$ and $\tilde{T}^{2}(\tilde{\beta}_{2}^{+})\in\tilde{J}$, again by property 3., so $\tilde{T}$ is stable in this case as well.
Because of this example, the following already appears to be a non-trivial question to answer:
Question 7.4.
@@
The boundary of a stable set is where the topology of $X$ changes, so it is natural to call it the bifurcation locus. We conjecture that is impossible for a bifurcation to an infinite type map to occur by moving from the interior of the stable region to its boundary:
Conjecture 7.6.
Maps in the boundary of a stable region are of finite type.
Recall that the return map to any interval of $X$ is equivalent to an $\mathrm{IET}$. In an upcoming work, we will prove the following theorem:
Theorem 7.7.
Any map contained in the boundary of a stable region has a critical connection. ∎
Because of Theorem 7.7, we get that any finite type map $T$ for which there is a return map to an interval of $X$ that satisfies the Keane condition (no critical connection for the induced $\mathrm{IET}$, see [MR2219821], [MR2648692]) and has at least $3$ branches (and thus does not satisfy Matching) is not in the closure of any stable region. For completeness, we give an example of such a map on $4$ intervals, which can be easily generalized to any dimension.
Figure 12: Dynamics of the map $T$ from Example 7.8.
Example 7.8 (Finite type map not in the closure of any stable component).
Consider the map $T$ with dynamics as shown in Figure 12. Its non-wandering set consists of the orbit of a single interval $J$ containing $\beta_{3}$ such that the return map $R_{J}$ to $J$ has three branches $J_{1},J_{2}$ and $J_{3}$, and the associated permutation is $\sigma$ is equal to $(321)$. Moreover, we assume that:
  1. 1.
    $R_{J}(J_{3})=T(J_{3})$;
  2. 2.
    $T(J_{1})\in I_{1},R_{J}(J_{1})=T^{2}(J_{1})$;
  3. 3.
    $T(J_{2})\in I_{2},R_{J}(J_{2})=T^{2}(J_{2})$;
  4. 4.
    $T^{2}(T)=T^{3}(I)=X$ (this can be achieved by choosing the sizes of the intervals $J_{1},J_{2},J_{3}$ and the parts outside of the orbit of $J$ accordingly, as in Figure 12);
  5. 5.
    The lengths $|J_{1}|=r_{1},|J_{2}|=r_{2},|J_{3}|=r_{3}$ of these three intervals are irrational numbers that are rationally independent, and $r_{3}>r_{1}$.
Thus $T$ is a finite type map that does not satisfy Matching. Moreover, by property $4$, we have that the translation factors associated to the three branches of the return map are $r_{2}+r_{3},r_{3}-r_{1},-r_{1}-r_{2}$, in that order. Thus the return map $R_{J}$ is equivalent to an $\mathrm{IET}$ that satisfies the Keane condition. Indeed, by rescaling we may assume that $J=I$, so our assumptions turn to: $r_{1}+r_{2}+r_{3}=1$, and $r_{1},r_{2},1$ are rationally independent. A critical connection would mean that there is a critical point that either lands on itself or on some other critical point. In the first case, we have that the following equation holds, where $n_{1},\dots,n_{3}\in\mathbb{N}$:
$n_{1}(1-r_{1})+n_{2}(1-2r_{1}-r_{2})+n_{3}(-r_{1}-r_{2}).$
Rearranging, we get that:
$(n_{1}-2n_{2}-n_{3})r_{1}+(-n_{2}-n_{3})r_{2}+(n_{1}+n_{2})=0,$
from which it easily follows that $n_{1}=n_{2}=n_{3}=0$. In the second case, we get that:
$r_{1}+n_{1}(1-r_{1})+n_{2}(1-2r_{1}-r_{2})+n_{3}(-r_{1}-r_{2})-r_{1}-r_{2}=0,$
where $n_{1},\dots,n_{3}\in\mathbb{Z}$. This is equivalent to:
$(n_{1}-2n_{2}-n_{3})r_{1}+(-n_{2}-n_{3}-1)r_{2}+(n_{1}+n_{2})=0,$
which is clearly impossible. Thus $T$ is not in the closure of any stable region.
Thus a necessary condition for a finite type map to be in the boundary of a stable region is that it has a critical connection. In Example 7.8, we can append new intervals to $I$ on which a critical connection happens, and then rescale the domain back to $I$. By the same reasoning as before, the resulting map is not in the boundary of any stable region, but it has a critical connection. Thus we see that having a critical connection is not a sufficient condition to be in the boundary of a stable region. A natural question to ask is whether the maximal possible number of critical connections is sufficient for a map to be in the closure of a stable region. The eventually periodic maps are exactly such maps, so we conjecture the following:
@@
Next, we turn to the topological properties of the set of all infinite type maps. Recall the following figure showing the set of infinite type maps for the special family from [MR2013352]:
This figure suggests the following conjecture:
Conjecture 7.10 (Accumulation of infinite type maps).
Arbitrarily close to any stable region, there is an infinite type map.
Finally, there is the question of the measure-theoretic analogue of the prevalence of finite type maps in the parameter space. Our Main Theorem II shows that the finite type maps are topologically prevalent in the parameter space since they contain an open and dense set. In [MR2013352], the authors show that the finite type maps in their family form a full measure subset of the parameter space. We conjecture that this is true in general as well:
Conjecture 7.11.
Infinite type maps form a measure zero subset of the full parameter space $\mathrm{ITM}(r)$.
A first step towards tackling both Conjecture 7.10 and Conjecture 7.11 would be to understand how the stable regions are arranged in $\mathrm{ITM}(r)$. Are they necessarily separated from each other, or can they ‘tile’ the parameter space?
Question 7.12.
Are the stable regions (partially) separated from each other? In other words, does for every stable region $[T]$ exist an open subset of the (outer) boundary on which the set of points that are contained in the boundary of some other stable component is small (e.g. meagre or countable)?

### v4

@@ 8.1 The Bruin–Troubetzkoy family
8.2 Future work: geometry and dynamics in the boundary of stable regions
In this final subsection, we discuss the properties of the full parameter space $\mathrm{ITM}(r)$ of all $\mathrm{ITM}$s on $r$ intervals, and we present a few conjectures. Recall that a connected component of the set of stable maps $\mathcal{S}(r)$ corresponds to the maximal neighbourhood $[T]$ of a stable map $T$ that satisfies the three properties from the definition of stability (Definition 5.1). Such sets $[T]$ are called stable regions. A very interesting problem is to describe the geometry of these regions.
Question 8.2.
@@ 8.2 Future work: geometry and dynamics in the boundary of stable regions
In a companion paper, we will give an example using ghost preimages (see Definition 5.8 and Example 5.6) which shows that stable regions are not convex in general. Since ghost preimages are codimension $1$ phenomena, it seems plausible that by taking the closure of a stable region, we can make it convex.
Question 8.3.
Is the closure of a stable region always convex?
Because the stable regions are generally not convex, the following also appears to be a non-trivial question:
Question 8.4.
@@
Because of ghost preimages, the boundary of stable regions can contain countably many hyperplanes. The accumulation properties of these hyperplanes seem difficult to describe, and we are led to the following question:
Question 8.5.
Is the boundary of every stable region locally connected?
Besides their geometry, a step towards the proof of Boshernitzan–Kornfeld Conjecture is to understand how the stable regions are arranged in the parameter $\mathrm{ITM}(r)$. Are they necessarily separated from each other, or can they ‘tile’ the parameter space?
Question 8.6.
Are the stable regions (partially) separated from each other? In other words, does for every stable region $[T]$ exist an open subset of the (outer) boundary on which the set of points that are contained in the boundary of some other stable component is small (e.g. meager or countable)?
@@
The boundary of a stable set is where the topology of a non-wandering set changes, so it is contained in bifurcation locus of $\mathrm{ITM}(r)$. Thus the answer to this question would constitute the start of the bifurcation theory of $\mathrm{ITM}$s. Intuitively, the bifurcations should be associated with the changes of critical itineraries, and they should in turn be associated with critical connections. This is confirmed by Theorem 7.5 for finite type maps. It is not clear that this should also hold for infinite-type maps:
Question 8.8.
Do infinite type maps in the boundary of a stable region have critical connections?
Since a necessary condition for a finite type map to be in the boundary of a stable region is that it has a critical connection, a natural question to ask is whether the maximal possible number of critical connections is sufficient for a map to be in the closure of a stable region. The eventually periodic maps are exactly such maps, so we conjecture the following: